(c) Let u = Re e +52+3+1. Show that u is harmonic function and find the harmonic conjugate v of u. [3]


differential equations
Answer #1

# U= Relezo+52² +22+1). Let z = titig 2² = t1²-yz + atlyé 2² = (²-yz > t ė (atly) i zl= (t²-ya? - atly? + Bryllt?- y2). 2²+52әo парамро В - -[41*-* )+ s- + 413 + 3а+ | (2 (41*-*) (341) + (21) ) - Xin ((+* - ) +1+5+ 3 ) . С ВАС - -5)+ +13 (21) --- elt²-y) ((til-ya +s) - 92y +30+ [ (Alican-yz) tlotila +41371²_42) +10)] trop(8ty(+2-ya) + loty+3y) ( Bylti-ye) tootly tlogI Now. consideo! za + sz²+ 3+1 = (Hi-yu)2 + schi-y?) – 4th?y? +36+1 S tel8ty (82-y?) tistry 3y) Now, Iv=img.(e29+523+3+11 ) I